OpenStudy (anonymous):

Which property does this? 2) 5x +15 +9 = 59 3) 5x + 24 = 59

OpenStudy (rina.r):

non

OpenStudy (snowfire):

Associative? 5x+(15+9)=59

OpenStudy (snowfire):

Just a guess though, at first glance I wouldn't have categorized it under a property.

Directrix (directrix):

To me, it looks like "definition of addition" which is not a property.

Directrix (directrix):

@Destinimichelle What was step 1 which was not posted?

OpenStudy (anonymous):

Yeah. My assignment says to list the 'reasons' for each statement and these are the statements: 1) 5(x+3) + 9 = 59 | The reason would be "Given" because that was the given problem. 2) 5x +15 + 9 = 59 | The reason would be Distributive property because you distribute the problem above to get that answer. but I can't figure out what 3 would be. 3) 5x +24 - 59 The options are: reflexive property transitive property addition property of equality multiplication property of equality subtraction property of equality division property of equality substitution property additive identity multiplicative property commutative property and associative property.

OpenStudy (anonymous):

3) 5x + 25 = 59 not - 59

OpenStudy (snowfire):

I think I'm sticking by my first answer of associative property of addition.

Directrix (directrix):

I might consider "substitution property" knowing that 24 can be substituted for (15 + 9). But, if substitution were the reason, then a step stating that 24 = 15 +9 should have been included. So, I'm back to not knowing.

OpenStudy (anonymous):

Thanks for the help.

Directrix (directrix):

@Destinimichelle Are you writing the steps or are they already given for you to justify?

OpenStudy (snowfire):

I think it is associative because generally the purpose of bringing up the associative property in a proof is to give one permission to combine some terms before you combine the rest. In this case, the 15 and 9 were chosen to be combined before anything else.

Directrix (directrix):

@Snowfire I agree with your rationale.

OpenStudy (anonymous):

Already given for me to justify.

OpenStudy (snowfire):

Well it seems my opinion is supported now :D did that make sense Destini?

OpenStudy (anonymous):

Yes, thank you! I asked my sister first for help and she was like "I'd go with Associate Property, but I haven't been in geometry for four years." So I thought I'd try here. I'm going with it since she said and you too. Thanks.

OpenStudy (snowfire):

Glad I could be of help, I think I just learned something new too

OpenStudy (anonymous):

Haha everyone's useful. Thanks again.

Directrix (directrix):

No step 4?

OpenStudy (snowfire):

Oh right, there should probably be another step if you ever want to get the answer ^^ maybe Destini already got it.

OpenStudy (anonymous):

Yeah. The rest of the steps are as follows: 4) 5x + 0 = 35 5) 5x = 35 60 x = 7

OpenStudy (snowfire):

Step 4 is probably addition property of equality, if memory serves me. 5 is additive identity, and 6 is division property of equality. I'm almost certain that's it, did you get the same?

Directrix (directrix):

On step 4, subtraction property of equality could be posited.

OpenStudy (snowfire):

It could go either way, I usually just think of the two properties as one and the same.

Directrix (directrix):

Me, too. @Snowfire

OpenStudy (anonymous):

I wasn't certain about 4, because my sister said additive identity, but at first I thought subtraction property of equality.

OpenStudy (snowfire):

The equality properties just state that the equation is unchanged if you add/subtract/multiply/divide by the same factor on both sides.

OpenStudy (anonymous):

Yeah. I know the equality properties state that.